$SPAD/src/input space3.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool space3.output
)set message test on
)set message auto off
)clear all

———-

Test for ThreeSpace domain — * —

--S 1 of 185
space := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (1)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 1
--S 2 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (2)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 2
--S 3 of 185
p1 := point [1,0,0]$(Point DoubleFloat)
--R
--R
--R   (3)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 3
--S 4 of 185
p2 := point [0,0,0]$(Point DoubleFloat)
--R
--R
--R   (4)  [0.,0.,0.]

--R                                                     Type: Point(DoubleFloat)
--E 4
--S 5 of 185
p3 := point [0,1,1]$(Point DoubleFloat)
--R
--R
--R   (5)  [0.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 5
--S 6 of 185
closedCurve(space,[p0,p1,p2,p3])
--R
--R
--R   (6)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 6
--S 7 of 185
closedCurve(space,[[1,1,1],[1,0,0],[0,0,0],[0,1,1]])
--R
--R
--R   (7)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 7
--S 8 of 185
cspace := closedCurve([p0,p1,p2,p3])$(ThreeSpace DoubleFloat)
--R
--R
--R   (8)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 8
--S 9 of 185
closedCurve cspace
--R
--R
--R   (9)  [[1.,1.,1.],[1.,0.,0.],[0.,0.,0.],[0.,1.,1.]]
--R                                               Type: List(Point(DoubleFloat))
--E 9
)clear all
--S 10 of 185
space := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (1)  3-Space with 0 components

--R                                                Type: ThreeSpace(DoubleFloat)
--E 10
--S 11 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (2)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 11
--S 12 of 185
p1 := point [1,0,0]$(Point DoubleFloat)
--R
--R
--R   (3)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 12
--S 13 of 185
p2 := point [0,0,0]$(Point DoubleFloat)
--R
--R
--R   (4)  [0.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 13
--S 14 of 185
p3 := point [0,1,1]$(Point DoubleFloat)
--R
--R
--R   (5)  [0.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 14
--S 15 of 185
closedCurve(space,[p0,p1,p2,p3])
--R
--R
--R   (6)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 15
--S 16 of 185
closedCurve? space
--R
--R
--R   (7)  true
--R                                                                Type: Boolean
--E 16

--S 17 of 185
curve(space,[p0,p1,p2,p3])
--R
--R
--R   (8)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 17
--S 18 of 185
point(space,p0)
--R
--R
--R   (9)  3-Space with 3 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 18
--S 19 of 185
components(space)
--R
--R
--R   (10)
--R   [3-Space with 1 component,3-Space with 1 component,3-Space with 1 component]
--R                                          Type: List(ThreeSpace(DoubleFloat))
--E 19
--S 20 of 185
space1 := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (11)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 20
--S 21 of 185
curve(space1,[p0,p1,p2,p3])
--R
--R
--R   (12)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 21
--S 22 of 185
point(space1,p0)
--R
--R
--R   (13)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 22

--S 23 of 185
space2 := point(p0)$(ThreeSpace DoubleFloat)
--R
--R
--R   (14)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 23
--S 24 of 185
space3 := curve[p0,p1,p2]$(ThreeSpace DoubleFloat)
--R
--R
--R   (15)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 24
--S 25 of 185
composite [space1,space2,space3]
--R
--R
--R   (16)  3-Space with 4 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 25
--S 26 of 185
curve(space,[p0,p1,p2,p3])
--R
--R
--R   (17)  3-Space with 4 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 26
--S 27 of 185
point(space,p0)
--R
--R
--R   (18)  3-Space with 5 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 27
--S 28 of 185
point(space,p1)
--R
--R
--R   (19)  3-Space with 6 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 28
--S 29 of 185
closedCurve(space,[p0,p1,p2])

--R
--R
--R   (20)  3-Space with 7 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 29
--S 30 of 185
composite [space1,space2,space3]
--R
--R
--R   (21)  3-Space with 4 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 30
--S 31 of 185
composites(space)
--R
--R
--R   (22)  []
--R                                          Type: List(ThreeSpace(DoubleFloat))
--E 31
--S 32 of 185
curve(space,[p0,p1,p2,p3])
--R
--R
--R   (23)  3-Space with 8 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 32
--S 33 of 185
point(space,p0)
--R
--R
--R   (24)  3-Space with 9 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 33
--S 34 of 185
space4 := copy space
--R
--R
--R   (25)  3-Space with 9 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 34
--S 35 of 185
curve(space,[p0,p1,p2])
--R
--R

--R   (26)  3-Space with 10 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 35
--S 36 of 185
point(space,p0)
--R
--R
--R   (27)  3-Space with 11 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 36
--S 37 of 185
sub := subspace(space)
--R
--R
--R   (28)  3-Space with depth of 3 and 11 components
--R                                                Type: SubSpace(3,DoubleFloat)
--E 37
--S 38 of 185
spNew := create3Space(sub)$(ThreeSpace DoubleFloat)
--R
--R
--R   (29)  3-Space with 11 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 38
--S 39 of 185
curve(space,[p0,p1,p2,p3])
--R
--R
--R   (30)  3-Space with 12 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 39
--S 40 of 185
curve(space,[[1,1,1],[1,0,0],[0,0,0],[0,1,1]])
--R
--R
--R   (31)  3-Space with 13 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 40
--S 41 of 185
cspace := curve([p0,p1,p2,p3])$(ThreeSpace DoubleFloat)
--R
--R
--R   (32)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)

--E 41
--S 42 of 185
curve cspace
--R
--R
--R   (33)  [[1.,1.,1.],[1.,0.,0.],[0.,0.,0.],[0.,1.,1.]]
--R                                               Type: List(Point(DoubleFloat))
--E 42
)clear all
--S 43 of 185
space := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (1)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 43
--S 44 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (2)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 44
--S 45 of 185
p1 := point [1,0,0]$(Point DoubleFloat)
--R
--R
--R   (3)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 45
--S 46 of 185
p2 := point [0,0,0]$(Point DoubleFloat)
--R
--R
--R   (4)  [0.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 46
--S 47 of 185
p3 := point [0,1,1]$(Point DoubleFloat)
--R
--R
--R   (5)  [0.,1.,1.]
--R                                                     Type: Point(DoubleFloat)

--E 47
--S 48 of 185
curve(space,[p0,p1,p2,p3])
--R
--R
--R   (6)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 48
--S 49 of 185
curve? space
--R
--R
--R   (7)  true
--R                                                                Type: Boolean
--E 49
)clear all
--S 50 of 185
space := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (1)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 50
--S 51 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (2)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 51
--S 52 of 185
p1 := point [1,0,0]$(Point DoubleFloat)
--R
--R
--R   (3)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 52
--S 53 of 185
p2 := point [0,0,0]$(Point DoubleFloat)
--R
--R
--R   (4)  [0.,0.,0.]
--R                                                     Type: Point(DoubleFloat)

--E 53
--S 54 of 185
p3 := point [0,1,1]$(Point DoubleFloat)
--R
--R
--R   (5)  [0.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 54
--S 55 of 185
closedCurve(space,[p0,p1,p2,p3])
--R
--R
--R   (6)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 55
--S 56 of 185
curve(space,[p0,p1,p2,p3])
--R
--R
--R   (7)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 56
--S 57 of 185
point(space,p0)
--R
--R
--R   (8)  3-Space with 3 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 57
--S 58 of 185
point(space,p3)
--R
--R
--R   (9)  3-Space with 4 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 58
--S 59 of 185
polygon(space,[p0,p1,p3])
--R
--R
--R   (10)  3-Space with 5 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 59

--S 60 of 185
polygon(space,[p0,p2,p3])
--R
--R
--R   (11)  3-Space with 6 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 60
--S 61 of 185
lllip(space)
--R
--R
--R   (12)  [[[1,2,3,4]],[[5,6,7,8]],[[9]],[[10]],[[11],[12,13]],[[14],[15,16]]]
--R                                   Type: List(List(List(NonNegativeInteger)))
--E 61
--S 62 of 185
closedCurve(space,[p0,p1,p2,p3])
--R
--R
--R   (13)  3-Space with 7 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 62
--S 63 of 185
curve(space,[p0,p1,p2,p3])
--R
--R
--R   (14)  3-Space with 8 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 63
--S 64 of 185
point(space,p0)
--R
--R
--R   (15)  3-Space with 9 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 64
--S 65 of 185
polygon(space,[p0,p1,p3])
--R
--R
--R   (16)  3-Space with 10 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 65
--S 66 of 185
llprop(space)

--R
--R
--R   (17)
--R   [[Component is closed, not solid], [Component is not closed, not solid],
--R    [Component is not closed, not solid], [Component is not closed, not solid],
--R    [Component is not closed, not solid,Component is not closed, not solid],
--R    [Component is not closed, not solid,Component is not closed, not solid],
--R    [Component is closed, not solid], [Component is not closed, not solid],
--R    [Component is not closed, not solid],
--R    [Component is not closed, not solid,Component is not closed, not solid]]
--R                                  Type: List(List(SubSpaceComponentProperty))
--E 66
--S 67 of 185
lprop(space)
--R
--R
--R   (18)
--R   [Component is not closed, not solid, Component is not closed, not solid,
--R    Component is not closed, not solid, Component is not closed, not solid,
--R    Component is not closed, not solid, Component is not closed, not solid,
--R    Component is not closed, not solid, Component is not closed, not solid,
--R    Component is not closed, not solid, Component is not closed, not solid]
--R                                        Type: List(SubSpaceComponentProperty)
--E 67
--S 68 of 185
closedCurve(space,[p0,p1,p2,p3])
--R
--R
--R   (19)  3-Space with 11 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 68
--S 69 of 185
curve(space,[p0,p1,p2,p3])
--R
--R
--R   (20)  3-Space with 12 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 69
--S 70 of 185
point(space,p0)
--R
--R
--R   (21)  3-Space with 13 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 70

--S 71 of 185
polygon(space,[p0,p1,p3])
--R
--R
--R   (22)  3-Space with 14 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 71
--S 72 of 185
lp(space)
--R
--R
--R   (23)
--R   [[1.,1.,1.], [1.,0.,0.], [0.,0.,0.], [0.,1.,1.], [1.,1.,1.], [1.,0.,0.],
--R    [0.,0.,0.], [0.,1.,1.], [1.,1.,1.], [0.,1.,1.], [1.,1.,1.], [1.,0.,0.],
--R    [0.,1.,1.], [1.,1.,1.], [0.,0.,0.], [0.,1.,1.], [1.,1.,1.], [1.,0.,0.],
--R    [0.,0.,0.], [0.,1.,1.], [1.,1.,1.], [1.,0.,0.], [0.,0.,0.], [0.,1.,1.],
--R    [1.,1.,1.], [1.,1.,1.], [1.,0.,0.], [0.,1.,1.], [1.,1.,1.], [1.,0.,0.],
--R    [0.,0.,0.], [0.,1.,1.], [1.,1.,1.], [1.,0.,0.], [0.,0.,0.], [0.,1.,1.],
--R    [1.,1.,1.], [1.,1.,1.], [1.,0.,0.], [0.,1.,1.]]
--R                                               Type: List(Point(DoubleFloat))
--E 72
--S 73 of 185
enterPointData(space,[p0,p1,p2,p3])
--R
--R
--R   (24)  44
--R                                                        Type: PositiveInteger
--E 73
)clear all
--S 74 of 185
space1 := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (1)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 74
--S 75 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (2)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 75
--S 76 of 185

p1 := point [1,0,0]$(Point DoubleFloat)
--R
--R
--R   (3)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 76
--S 77 of 185
p2 := point [0,0,0]$(Point DoubleFloat)
--R
--R
--R   (4)  [0.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 77
--S 78 of 185
p3 := point [0,1,1]$(Point DoubleFloat)
--R
--R
--R   (5)  [0.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 78
--S 79 of 185
curve(space1,[p0,p1,p2,p3])
--R
--R
--R   (6)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 79
--S 80 of 185
space2 := copy space1
--R
--R
--R   (7)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 80
--S 81 of 185
point(space1,p3)
--R
--R
--R   (8)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 81
--S 82 of 185
space3 := copy space1
--R

--R
--R   (9)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 82
--S 83 of 185
curve(space3,[p0,p1,p2])
--R
--R
--R   (10)  3-Space with 3 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 83
--S 84 of 185
newSpace1 := merge [space1,space2,space3]
--R
--R
--R   (11)  3-Space with 6 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 84
--S 85 of 185
newSpace2 := merge(space2,space3)
--R
--R
--R   (12)  3-Space with 4 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 85
--S 86 of 185
space := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (13)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 86
--S 87 of 185
prop := new()$SubSpaceComponentProperty()
--R
--R
--R   (14)  Component is not closed, not solid
--R                                              Type: SubSpaceComponentProperty
--E 87
--S 88 of 185
lprop := [prop, prop, prop]
--R
--R
--R   (15)

--R   [Component is not closed, not solid, Component is not closed, not solid,
--R    Component is not closed, not solid]
--R                                        Type: List(SubSpaceComponentProperty)
--E 88
--S 89 of 185
mesh(space,[[p0,p1,p2],[p1,p2,p3],[p0,p2,p3]],lprop,prop)
--R
--R
--R   (16)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 89
--S 90 of 185
mesh(space,[[p0,p1,p2],[p1,p2,p3],[p0,p2,p3]],lprop,prop)
--R
--R
--R   (17)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 90
--S 91 of 185
mesh(space,[[p0,p1,p2],[p1,p2,p3],[p0,p2,p3]],closed?(prop),closed?(prop))
--R
--R
--R   (18)  3-Space with 3 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 91
--S 92 of 185
b := close(prop,true)
--R
--R
--R   (19)  true
--R                                                                Type: Boolean
--E 92
--S 93 of 185
mesh(space,[[p0,p1,p2],[p1,p2,p3],[p0,p2,p3]],b,b)
--R
--R
--R   (20)  3-Space with 4 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 93
--S 94 of 185
mesh(space,[[p0,p1,p2],[p1,p2,p3],[p0,p2,p3]],closed?(prop),closed?(prop))
--R
--R
--R   (21)  3-Space with 5 components

--S 95 of 185
mesh(space,[[[1,1,1],[1,0,0],[0,0,0]],_
            [[1,0,0],[0,0,0],[0,1,1]],_
            [[1,1,1],[0,0,0],[0,1,1]]],closed?(prop),closed?(prop))
--R
--R
--R   (22)  3-Space with 6 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 95
--S 96 of 185
mesh(space,[[[1,1,1],[1,0,0],[0,0,0]],_
            [[1,0,0],[0,0,0],[0,1,1]],_
            [[1,1,1],[0,0,0],[0,1,1]]],b,b)
--R
--R
--R   (23)  3-Space with 7 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 96
)clear all
--S 97 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (1)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 97
--S 98 of 185
p1 := point [1,0,0]$(Point DoubleFloat)
--R
--R
--R   (2)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 98
--S 99 of 185
p2 := point [0,0,0]$(Point DoubleFloat)
--R
--R
--R   (3)  [0.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 99
--S 100 of 185

p3 := point [0,1,1]$(Point DoubleFloat)
--R
--R
--R   (4)  [0.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 100
--S 101 of 185
space := mesh([[p0,p1,p2],[p1,p2,p3],[p0,p2,p3]])$(ThreeSpace DoubleFloat)
--R
--R
--R   (5)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 101
--S 102 of 185
space1 := mesh([[p0,p1,p2],_
                [p1,p2,p3],_
                [p0,p2,p3]],false,false)$(ThreeSpace DoubleFloat)
--R
--R
--R   (6)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 102
)clear all
--S 103 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (1)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 103
--S 104 of 185
p1 := point [1,0,0]$(Point DoubleFloat)
--R
--R
--R   (2)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 104
--S 105 of 185
p2 := point [0,0,0]$(Point DoubleFloat)
--R
--R
--R   (3)  [0.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 105

--S 106 of 185
p3 := point [0,1,1]$(Point DoubleFloat)
--R
--R
--R   (4)  [0.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 106
--S 107 of 185
space := mesh([[p0,p1,p2],[p1,p2,p3],[p0,p2,p3]])$(ThreeSpace DoubleFloat)
--R
--R
--R   (5)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 107
--S 108 of 185
mesh(space)
--R
--R
--R   (6)
--R   [[[1.,1.,1.],[0.,0.,0.],[0.,1.,1.]], [[1.,0.,0.],[0.,0.,0.],[0.,1.,1.]],
--R    [[1.,1.,1.],[1.,0.,0.],[0.,0.,0.]]]
--R                                         Type: List(List(Point(DoubleFloat)))
--E 108
--S 109 of 185
s := mesh([[p0,p1,p2],[p1,p2,p3],[p0,p2,p3]])$(ThreeSpace DoubleFloat)
--R
--R
--R   (7)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 109
--S 110 of 185
mesh(s)
--R
--R
--R   (8)
--R   [[[1.,1.,1.],[0.,0.,0.],[0.,1.,1.]], [[1.,0.,0.],[0.,0.,0.],[0.,1.,1.]],
--R    [[1.,1.,1.],[1.,0.,0.],[0.,0.,0.]]]
--R                                         Type: List(List(Point(DoubleFloat)))
--E 110
--S 111 of 185
space2 := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (9)  3-Space with 0 components

--S 112 of 185
curve(space2,[p0,p1,p2,p3])
--R
--R
--R   (10)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 112
--S 113 of 185
mesh?(space2)
--R
--R
--R   (11)  false
--R                                                                Type: Boolean
--E 113
--S 114 of 185
s1 := mesh([[p0,p1,p2],[p1,p2,p3],[p0,p2,p3]])$(ThreeSpace DoubleFloat)
--R
--R
--R   (12)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 114
--S 115 of 185
mesh?(s1)
--R
--R
--R   (13)  true
--R                                                                Type: Boolean
--E 115
--S 116 of 185
i := enterPointData(space2,[p0,p1,p2,p3])::NNI
--R
--R
--R   (14)  8
--R                                                     Type: NonNegativeInteger
--E 116
--S 117 of 185
modifyPointData(space2,i,p2)
--R
--R
--R   (15)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 117

--S 118 of 185
point(space2,p0)
--R
--R
--R   (16)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 118
--S 119 of 185
curve(space2,[p0,p1,p2,p3])
--R
--R
--R   (17)  3-Space with 3 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 119
--S 120 of 185
numberOfComponents(space2)
--R
--R
--R   (18)  3
--R                                                        Type: PositiveInteger
--E 120
)clear all
--S 121 of 185
space1 := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (1)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 121
--S 122 of 185
numberOfComposites(space1)
--R
--R
--R   (2)  0
--R                                                     Type: NonNegativeInteger
--E 122
--S 123 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (3)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 123

--S 124 of 185
p1 := point [1,0,0]$(Point DoubleFloat)
--R
--R
--R   (4)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 124
--S 125 of 185
p2 := point [0,0,0]$(Point DoubleFloat)
--R
--R
--R   (5)  [0.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 125
--S 126 of 185
p3 := point [0,1,1]$(Point DoubleFloat)
--R
--R
--R   (6)  [0.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 126
--S 127 of 185
curve(space1,[p0,p1,p2,p3])
--R
--R
--R   (7)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 127
--S 128 of 185
point(space1,p0)
--R
--R
--R   (8)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 128
--S 129 of 185
space2 := point(p0)$(ThreeSpace DoubleFloat)
--R
--R
--R   (9)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 129
--S 130 of 185

space3 := curve [p0,p1,p2]$(ThreeSpace DoubleFloat)
--R
--R
--R   (10)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 130
--S 131 of 185
s := composite [space1,space2,space3]
--R
--R
--R   (11)  3-Space with 4 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 131
--S 132 of 185
numberOfComposites(s)
--R
--R
--R   (12)  1
--R                                                        Type: PositiveInteger
--E 132
--S 133 of 185
space := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (13)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 133
--S 134 of 185
point(space,p0)
--R
--R
--R   (14)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 134
--S 135 of 185
curve(space,[p0,p1,p2,p3])
--R
--R
--R   (15)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 135
--S 136 of 185
closedCurve(space,[p0,p1,p2])
--R

--R
--R   (16)  3-Space with 3 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 136
--S 137 of 185
objects space
--R
--R
--R   (17)  [points= 1,curves= 2,polygons= 0,constructs= 0]
--RType: Record(points: NonNegativeInteger,curves: NonNegativeInteger,polygons: NonNegativeInteger,constructs: NonNegativeInteger)
--E 137
)clear all
--S 138 of 185
s := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (1)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 138
--S 139 of 185
p := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (2)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 139
--S 140 of 185
point(s,p)
--R
--R
--R   (3)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 140
--S 141 of 185
point(s,[1,1,1])
--R
--R
--R   (4)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 141
--S 142 of 185
p0 := point [1,0,0]$(Point DoubleFloat)
--R

--R
--R   (5)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 142
--S 143 of 185
point(s,p)
--R
--R
--R   (6)  3-Space with 3 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 143
--S 144 of 185
i := enterPointData(s,[p0])::NNI
--R
--R
--R   (7)  4
--R                                                     Type: NonNegativeInteger
--E 144
--S 145 of 185
point(s,i)
--R
--R
--R   (8)  3-Space with 4 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 145
--S 146 of 185
p := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (9)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 146
--S 147 of 185
space := point(p)$(ThreeSpace DoubleFloat)
--R
--R
--R   (10)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 147
)clear all
--S 148 of 185
s := create3Space()$(ThreeSpace DoubleFloat)
--R

--R
--R   (1)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 148
--S 149 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (2)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 149
--S 150 of 185
curve(s,[p0,p0])
--R
--R
--R   (3)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 150
--S 151 of 185
space1 := point(p0)$(ThreeSpace DoubleFloat)
--R
--R
--R   (4)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 151
--S 152 of 185
point(space1)
--R
--R
--R   (5)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 152
)clear all
--S 153 of 185
s := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (1)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 153
--S 154 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R

--R
--R   (2)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 154
--S 155 of 185
curve(s,[p0,p0,p0])
--R
--R
--R   (3)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 155
--S 156 of 185
point? s
--R
--R
--R   (4)  false
--R                                                                Type: Boolean
--E 156
--S 157 of 185
space := point(p0)$(ThreeSpace DoubleFloat)
--R
--R
--R   (5)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 157
--S 158 of 185
point? space
--R
--R
--R   (6)  true
--R                                                                Type: Boolean
--E 158
)clear all
--S 159 of 185
space := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (1)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 159
--S 160 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R

--R
--R   (2)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 160
--S 161 of 185
p1 := point [1,0,0]$(Point DoubleFloat)
--R
--R
--R   (3)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 161
--S 162 of 185
p2 := point [0,0,0]$(Point DoubleFloat)
--R
--R
--R   (4)  [0.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 162
--S 163 of 185
p3 := point [0,1,1]$(Point DoubleFloat)
--R
--R
--R   (5)  [0.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 163
--S 164 of 185
polygon(space,[p0,p1,p2,p3])
--R
--R
--R   (6)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 164
--S 165 of 185
polygon(space,[[1,1,1],[0,0,-1],[1,0,1]])
--R
--R
--R   (7)  3-Space with 2 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 165
--S 166 of 185
s := polygon([p0,p1,p2,p3])$(ThreeSpace DoubleFloat)
--R
--R
--R   (8)  3-Space with 1 component

)clear all
--S 167 of 185
space := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (1)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 167
--S 168 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (2)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 168
--S 169 of 185
p1 := point [1,0,0]$(Point DoubleFloat)
--R
--R
--R   (3)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 169
--S 170 of 185
p2 := point [0,0,0]$(Point DoubleFloat)
--R
--R
--R   (4)  [0.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 170
--S 171 of 185
p3 := point [0,1,1]$(Point DoubleFloat)
--R
--R
--R   (5)  [0.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 171
--S 172 of 185
curve(space,[p0,p1,p2,p3])
--R
--R
--R   (6)  3-Space with 1 component

--S 173 of 185
s := polygon([p0,p1,p2,p3])$(ThreeSpace DoubleFloat)
--R
--R
--R   (7)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 173
--S 174 of 185
polygon s
--R
--R
--R   (8)  [[1.,0.,0.],[0.,0.,0.],[0.,1.,1.]]
--R                                               Type: List(Point(DoubleFloat))
--E 174
)clear all
--S 175 of 185
space := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (1)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 175
--S 176 of 185
p0 := point [1,1,1]$(Point DoubleFloat)
--R
--R
--R   (2)  [1.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 176
--S 177 of 185
p1 := point [1,0,0]$(Point DoubleFloat)
--R
--R
--R   (3)  [1.,0.,0.]
--R                                                     Type: Point(DoubleFloat)
--E 177
--S 178 of 185
p2 := point [0,0,0]$(Point DoubleFloat)
--R
--R
--R   (4)  [0.,0.,0.]

--S 179 of 185
p3 := point [0,1,1]$(Point DoubleFloat)
--R
--R
--R   (5)  [0.,1.,1.]
--R                                                     Type: Point(DoubleFloat)
--E 179
--S 180 of 185
curve(space,[p0,p1,p2,p3])
--R
--R
--R   (6)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 180
--S 181 of 185
polygon? space
--R
--R
--R   (7)  false
--R                                                                Type: Boolean
--E 181
--S 182 of 185
s := polygon([p0,p1,p2,p3])$(ThreeSpace DoubleFloat)
--R
--R
--R   (8)  3-Space with 1 component
--R                                                Type: ThreeSpace(DoubleFloat)
--E 182
--S 183 of 185
polygon s
--R
--R
--R   (9)  [[1.,0.,0.],[0.,0.,0.],[0.,1.,1.]]
--R                                               Type: List(Point(DoubleFloat))
--E 183
--S 184 of 185
space1 := create3Space()$(ThreeSpace DoubleFloat)
--R
--R
--R   (10)  3-Space with 0 components
--R                                                Type: ThreeSpace(DoubleFloat)
--E 184

--S 185 of 185
sub := subspace(space1)
--R
--R
--R   (11)  3-Space with depth of 3 and 0 components
--R                                                Type: SubSpace(3,DoubleFloat)
--E 185
)spool
)lisp (bye)

———-

References^{[1] nothing}

35